i need help with this question
suppose that the length of of a rectangle is 3 inches longer than the width and that the perimeter of the rectangle is 78.
set up an equation involving only L, the length of the rectangle.then solve the equation algebraically to find the length of the rectangle. find the width as well.
please show how you got the answer.







-

Answer by checkley71(8403)

L=W+3 ORE W=L-3
PERIMETER=2L+2W
78=2L+2(L-3)
78=2L+2L-6
78=4L-6
4L=78+6
4L=84
L=84/4
L=21 INCHES.
W=L-3
W=21-3
W=18 INCHES FOR THE WIDTH.
PROOF
2*21+2*18=78
42+36=78
78=78